• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R.A. Lindgren

    Log entry time 16:02:17 on January30,2003

    Entry number 93879

    keyword=shift summary

    Day Shift begins
    Compton polarimeter DAQ crashed during OWL shift, Restarted at 8:30 am by JP
    Continuing Runs 20627/1627/1568
    stopped runs 20627/1627/1568 at 8:30 am
    started runs 20628/1628/1569 Junk
    Started runs 20629/1629/1570 at 8:35 AM
    Stop runs 20629/1629/1570 at 8:59 AM
    MCC adjusting source to get more QE
    Restart Run
    Start Runs 20630/1630/1571 at 9:00 AM
    Stop Runs 20630/1630/1571 at 9:30 AM
    Finished 5.5.5 kinematics 34K good events total
    MCC repairing ARC,
    Mark Stephens request Hall A access
    Back to beam permit
    Page Saha to make energy measurement when beam returns
    Cates making polarization measurement 35.8%
    Want 5 ua, no raster, no chicane
    Waiting for beam tune.
    12:15 Pulsed beam in Hall. okay
    12:30 Tune CW beam in Hall
    13:00 MCC reports an FSD from the compton polarimeter
    End of run failed in Compton polarimeter
    Sirish advises to put electron detector in garage
    Can't remove from beam position. Error keeps popping up
    Reboot Compton polarimeter slow controls: It doesn't help
    13:15 Sirish on the way in.
    14:20 The electron polarimeter detector is stuck in the beam line due to some hardware problem, which requires an access to repair. So in lieu of an access we and the run coordinator decided to forgo the energy measurement and continue to take data. MCC is tuning the beam through the chicane and Sirish is monitoring using the polarimeter. We expect to be back taking data shortly

    15:00 Beam is tuned. Mask and raster beam. Stop the beam.
    Put target in. Measure polarization 38.3 %

    15:20 Start runs 20631/1631/1572
    16:00 Shift change Continue the same run.